OpenStudy (anonymous):

Which symbol creates a true sentence when x equals 7? 52 + (x – 3)2 __ 35 A. = B. > C. <

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

is it \[52+(x-3)^2~~~35\]?

OpenStudy (anonymous):

yeah i think so

OpenStudy (misty1212):

lol if you do not know, who does?

OpenStudy (anonymous):

@misty1212 can you help me plz

OpenStudy (misty1212):

sure but i have to know what the question is first you did not write exponents, i kinda had to guess

OpenStudy (misty1212):

so maybe it is \[52+(x-3)^2~~~35\] or maybe \[5^2+(x-3)^2~~~35\]

OpenStudy (anonymous):

52 + (x – 3)2 __ 35

OpenStudy (anonymous):

it's the 2 one

OpenStudy (misty1212):

oooh ok

OpenStudy (misty1212):

replace \(x\) by \(7\) to get \[5^2+(7-3)^2\] and now lets compute that

OpenStudy (misty1212):

what is \(5^2\)?

OpenStudy (misty1212):

if you don't know i will tell you

OpenStudy (jameshorton):

what the question

OpenStudy (anonymous):

yeah i dont know

OpenStudy (misty1212):

\[5^2=5\times 5=25\]

OpenStudy (misty1212):

and \(7-3=4\) so \((7-3)^2=4^2=16\)

OpenStudy (misty1212):

when you add you get \[5^2+4^2=25+16=41\] so it is bigger than 35

OpenStudy (misty1212):

go with B \(>\)

OpenStudy (anonymous):

`yes

OpenStudy (anonymous):

thx

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]